

********************************************************************
UARS WEEKLY STATUS REPORT - 28 SEP 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE has been in standby mode this week and will begin normal
operations again on October 1st.
***********
SUSIM
***********
SUSIM was turned off September 24. Two days later on September
26, the solar F10.7 radio flux reached its highest value (282.6)
of solar cycle 23. We are disappointed not to have measurements of
the solar UV irradiance on that day. Ironically, we continue to
receive science data files from the CDHF for the days we are
turned off.
When it was on, SUSIM completed its daily solar scans during
periods of predicted UARS telemetry.
We await the PACOR-A level 0 SUSIMA science files for evaluation.
After some detailed evaluation, the we believe that the level 0
SUSIM data can be processed using the RSS system on the SUSIM RAC,
as long as all of the level 0 files are present and are in the same
CDHF format.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
***** ALL PEM COMMUNICATION SHOULD NOW BE THROUGH THE PEM2K:: NODE *****
The PEM instrument is off at the current time.
At the present time, PEM energy deposition from its spectrometers is not
being generated. Recent data configuration changes have affected PEM software
causing multiple failures. These failures are being reviewed by the PEM team.
In order to correct for new telemetry requirements, PEM will need to update
processing code with a patch delivery. The delivery patch may include changes
which would allow the PEM software to execute on a DEC Alpha machine.
Currently, PEM has requested GSFC assistance with compiling for the DEC Alpha
since the GSFC executables run on the Alpha and SwRI's compiled versions do
not. It is unknown what the reprocessing state is since the CDHF may go away.
Updated HEPS calibration parameters have been finalized and are ready for
delivery. Updated MEPS quality calibration files have been finalized and are
ready for delivery. Level 3 software errors, discovered by Dr. Dean Pesnell,
have been fixed; his suggestions have been incorporated with in the PEM
production code. Both updated HEPS and MEPS calibration data, along with the
Level 3 code changes, will influence the in-situ energy deposition predictions,
so new versions of PEM energy deposition Level 3 files will be generated.
*Beginning of power orbit day (PEM activation): array current >5 amps, array
voltage >53.5 volts, and battery current >2.0 amps.
End of power orbit day (PEM deactivation): time in power orbit day >4 minutes,
array current <9 amps, and battery currents <0.2 amps.
Power orbit day variation: ~39-40 min at yaw to ~50-51 min at maximum beta.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
The MPG attended the weekly Technical Team Meeting
held on 27 September.
During this past week all UARS instruments, except
ACRIM, have been in their standby mode. (ACRIM has
remained on to avoid instrument power supply problems.)
After the Technical Team Meeting on 27 September, it
was decided to resume data collection by HALOE, SOLSTICE,
and SUSIM for one week, starting on 1 October 2001.
Instrument scheduling after 7 October remains to be
determined.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 15 September 2001 (Orbit 54736, GMT Day 258) through 21
September 2001 (Orbit 54840, GMT Day 264).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
HALOE 260/0013 54766 HALOE Power ON: Part-1
HALOE 260/0524 54769 HALOE Power ON: Part-2
The beta angle ranged between the angles of 14.4 degrees to 36.8 degrees for
this report period. The beta angle is now increasing towards a maximum angle
of 54 degrees on 30 September 2001 (DOY 273).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.0 to 1.2 degrees C. Additional battery performance data for each
battery on 14 September 2001 and 21 September 2001 are provided in Appendix D.
The clock error ranged from -9.1 to 13.1 msec during this report period. Clock
rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 00
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 21
¥ HALOE = 02
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ NBTR = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/243 - 2287.499091
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/256 show the observatory data loss to be 303 hours, 11
minutes, and 12.704 seconds. This shows a data loss of 25.680 seconds since the
last report period. This is a 0.39 percent data loss, which equals a 99.61
percent data capture for the mission.
The increase of 25.680 seconds occurred over two (2) days:
· 13.744 seconds on Day 259 due to a Data Hit
· 13.840 seconds on Day 260 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The FOT has begun the process of implementing Raytheon's Eclipse ground system
and continues to work on the TAC replacement system. These new systems will
provide enhanced capability and make up for the loss of some UARS facilities,
which have been shut down due to budgetary reasons. This will be a huge effort
and must be completed as soon as possible in order to provide a low cost method
for flying UARS. The FOT will need assistance from all entities involved to
insure the success of this operation. The estimated time for completion is
April 2002.
MEETINGS
The UARS FOT attended a Flight Software CCB on Tuesday, September 18, 2001.
The UARS FOT attended TAC RS Meeting on Tuesday, September 18, 2001.
The UARS FOT attended UARS Technical Team Meeting on Thursday, September 20,
2001.
The UARS FOT attended PACOR A Meeting on Friday, September 21, 2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 17 October 2001 (DOY 290). This
maneuver will be a Forward to Reverse Flight maneuver (Northern Hemisphere
viewing).
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 21 September 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
54749 258/2130 01-573 OBC/CLK Clk Adj -20.6 to 0.00 msec/day
54753 259/0412 01-574 OBC/CLK Clk Adj 0.00 to 20.6 msec/day
54763 259/1944 01-575 NBTR NOOPS D-260
54766 260/0013 01-576 HALOE Power ON: Part-1
54767 260/0226 01-578 OBC/CLK Clk Adj 20.6 to 0.00 msec/day
54769 260/0524 01-575 NBTR NBTR PBK & RTS-58
54769 260/0524 01-578 HALOE Power ON: Part-2
54772 260/0926 01-575 NBTR NOOPS D-260
54778 260/1905 01-579 NBTR NOOPS D-261
54784 261/0523 01-579 NBTR NOOPS D-261
54786 261/0847 01-579 NBTR NBTR PBK & RTS-58
54786 261/0847 01-580 NBTR NBTR STOP & REWIND
54793 261/1909 01-581 NBTR NOOPS D-262
54799 262/0525 01-581 NBTR PBK & RTS-58
54800 262/0604 01-581 NBTR NOOPS D-262
54804 262/1250 01-581 NBTR PBK & RTS-58
54808 262/1909 01-582 NBTR NOOPS D-263
54814 263/0525 01-582 NBTR PBK & RTS-58
54815 263/0605 01-582 NBTR NOOPS D-263
54823 263/1905 01-583 NBTR NOOPS D-264
54830 264/0602 01-583 NBTR Noops
54837 264/1724 01-583 NBTR NOOPS D264
54838 264/1904 01-584 NBTR NOOPS D265
________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 21 September 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
-None this report period
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 21 September 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
54749 258/2134 Clk Adj -20.6 to 0.00 msec/day
54753 259/0417 Clk Adj 0.00 to 20.6 msec/day
54767 260/0230 Clk Adj 20.6 to 0.00 msec/day
________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 21 September 2001
14 September 2001 - GMT Day 257 Beta = 10.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/ AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-78.4 +4.51/+3.87 +28.8/-13.6 26.1 1.030 19.2
3 +22.4/-16.8 +3.23/+2.60 +28.0/-13.2 25.8 1.030 18.8
21 September 2001 - GMT Day 264 Beta = 36.8 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/+0.0 11.8 0.000 0.0
2 +00.0/-190.4 +5.47/+4.51 +28.0/-14.4 25.1 1.030 19.2
3 +22.4/-44.8 +4.51/+3.23 +27.2/-14.8 25.1 1.030 18.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 21 September 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-057 NBTR 54786 261/084 N/A NBTR-B Unscheduled stop in PBK
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-057 NBTR 54786 261/0847 N/A NBTR-B Unscheduled stop in PBK
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-054 FOT see list Sept. 01 14170 Generic Late Acquisition
SEVNTFW01/38